Sam and pam's ages are consecutive integers. Three years ago the sum of there age was 15. How old are they now, with sam being older?
Found 2 solutions by josgarithmetic, Alan3354:
Answer by josgarithmetic(39618)

z and p ("Zam")

z=p+1 Sam is older.

(z-3)+(p-3)=15.

Order of z and p is not critical without other information.
z+p-6=15
z+p=21
-
If sam is older, then p+1+p=21
2p+1=21
p=10 and then z=11.



Answer by Alan3354(69443)

Sam and pam's ages are consecutive integers. Three years ago the sum of there age was 15. How old are they now, with sam being older?
=========
They will always be 1 year apart, consecutive integers.
15 = 7 + 8